# {textblock*}

### v1

@@ IV Numerical results
To accurately account for the subtleties in the model, we carry out numerical computations of the perturbation evolution based on the Transport method [56]. We consider three sets of benchmark parameters, as listed in Table 1: Case 1 corresponds to ultra-light PBHs, Case 2 to asteroid PBHs within the dark matter window, and Case 3 to PBHs detectable via microlensing. All three parameter sets yield slightly larger spectral indexes at the CMB pivot scale of $0.05\mathrm{Mpc}^{-1}$ , which remain consistent with the P-ACT results.
However, according to P-ACT-LB constraint on $\alpha_{s}$, Case 3 is excluded at more than the $2\sigma$ level. In fact, in our model, the production of PBHs with masses $\gtrsim 10^{20}\mathrm{g}$ is generally ruled out by ACT, as it is associated with a relatively large running of the spectral index (see Fig. 2). As a result, we conclude that PBHs larger than $\sim 10^{20}$ g are excluded as a dark matter candidate at 2 $\sigma$ level according to the P-ACT results.
Case 1 can be both consistent with the data but also potentially interesting. The ultra-light PBHs produced are expected to evaporate before Big Bang Nucleosynthesis (BBN) due to Hawking radiation [57, 58, 59]. Those evaporating tiny PBHs can help to reheat the universe, produce high frequency gravitational waves by Hawking radiation, and induce second order gravitational waves through the poltergeist mechanism[60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. From quantum gravity point of view, those evaporated PBHs may abundantly leave remnants as dark matter [72, 73, 74, 75].
In the following, we focus on the most compelling case, Case 2, where the near monochromatic asteroid mass PBHs produced from the model can constitute the whole dark matter [76, 77, 78], which shows great compatibility with ACT results.
@@
It is also worth mentioning that, as illustrated in Fig. 3, the enhancement of the small-scale power spectrum not only affects the large-scale spectral index $n_{s}$ and its running $\alpha_{s}$, which can be probed through direct measurements of the primordial power spectrum from future CMB surveys [79, 80], but also induces significant $\mu$-distortion [81, 82]—potentially detectable by PIXIE [25, 26] [^7] See detailed analysis in [11].. Additionally, the associated induced gravitational waves fall within the sensitivity ranges of future detectors such as LISA [27], DECIGO [84], Taiji [85]/TianQin [86] and BBO [55]. Therefore, upcoming missions could provide crucial observational evidence in support of this model.
Figure 4: The $n_{s}-r$ relation in the $R^{2}$ model and $R^{2}$-Higgs model together with the joint constraints from BICEP/Keck observations [87] (denoted as BK18) with Planck-LB (orange), ACT-LB (blue), and P-ACT-LB (purple). The constraint contours are taken from [6]. The dark blue dashed line with hollow dots refers to the $n_{s}-r$ relation in the $R^{2}$ model, and the light cyan dashed line and dots refer to the $R^{2}$-Higgs model with PBH production. For the light cyan plot, we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), $0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (star-shaped, Case 2), $0.0685$.
Figure 5: The $n_{s}-\alpha_{s}$ relation in $R^{2}$-Higgs model together with the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). The light cyan dots refer to the $\alpha_{s}-n_{s}$ relation in $R^{2}$-Higgs model with PBH production, for which we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), $0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (star-shaped, Case 2), $0.0685$.
@@ V Discussion and Conclusion
The recently released ACT observation joint with Planck data reports a slightly larger spectral index $n_{s}$ and a positive running $\alpha$ placing well-motivated $R^{2}$ inflation and Higgs inflation models under tension at $2\sigma$ level. In this letter, we have shown that we can resolve the tension within the simple combination of the two models, namely, introducing a non-minimally coupled scalar field to the $R^{2}$ model. By tuning the non-minimally coupling constant $\xi$, the model realizes a hybrid-like inflationary scenario, in which the non-minimally coupled scalar field contributes a blue-tilted component to the primordial power spectrum, yielding an increase in the $n_{s}$ and a positive $\alpha_{s}$ consistent with the current observation.
At the same time, the blue-tilted spectrum significantly enhances the fluctuation power at small scales, potentially leading to the formation of primordial black holes (PBHs). In Fig. 2, we show the relation between $n_{s}$, $\alpha_{s}$, and the mass of the resulting PBHs. To remain consistent with ACT results, We demonstrated that the observed values of $n_{s}$ and positive $\alpha_{s}$ may serve as indirect evidence of such small-scale enhancement, consistent with the formation of PBHs with masses $\lesssim 10^{20}~{}\mathrm{g}$, which may account for the entirety of dark matter. It is worth mentioning that we do not claim that the tension between the new data and $R^{2}$ model predictions must be resolved by PBHs that explain all dark matter, but rather demonstrate our model naturally aligns the modifications of large scale power spectrum with testable small-scale features in the model. More specifically, the small-scale power spectrum of our model can be tested by upcoming observations of $\mu$-distortion (e.g. via PIXIE) and the stochastic gravitational wave background (e.g. via LISA), offering promising avenues for observational confirmation.
@@ Acknowledgements.
We thank Misao Sasaki, Tsutomu Yanagida and Ying-li Zhang for useful discussions. This work is supported in part by JSPS KAKENHI Grant Nos. JP20H05853, JP24K00624 [X.W.]; JP23KF0289, JP24H01825, JP24K07027 [K.K.], and by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo [X.W.]. Kavli IPMU is supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. X.W. would also like to acknowledge the insightful input from the Random Discussion, held weekly at Kavli IPMU within the theoretical cosmology group.

### v2

@@ IV Numerical results
To account for the subtleties in the model, we perform numerical computations using the Transport method [72]. We consider three sets of benchmark parameters, as listed in Table 1: Case 1 corresponds to ultra-light PBHs that evaporate before Big Bang Nucleosynthesis (BBN), Case 2 to asteroid mass PBHs within the dark matter window, and Case 3 to PBHs potentially detectable through microlensing surveys. All three parameter sets yield a slightly larger spectral index at the CMB pivot scale of $0.05\mathrm{Mpc}^{-1}$, which remains consistent with the P-ACT results.
However, Case 3 is excluded at more than the $2\sigma$ level based on the P-ACT-LB constraint on $\alpha_{s}$. This is due to the fact that, in our model, the production of PBHs with masses $\gtrsim 10^{20}\mathrm{g}$ is generally disfavored by ACT, as it is associated with a relatively large running of the spectral index (see Fig. 3). As a result, we conclude that PBHs larger than $\sim 10^{20}$ g are excluded as a dark matter candidate in our model at 2 $\sigma$ level according to the P-ACT results.
Case 1 is not only consistent with current data but also phenomenologically intriguing. The ultra-light PBHs produced in this case are expected to evaporate before BBN via Hawking radiation [73, 74, 75]. Those evaporating tiny PBHs can help to reheat the universe, produce high-frequency gravitational waves by Hawking radiation, and induce second-order gravitational waves through the poltergeist mechanism [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]. From quantum gravity point of view, those evaporated PBHs may abundantly leave remnants as dark matter [88, 89, 90, 91].
In the following, we focus on the most compelling case, Case 2, where the near monochromatic asteroid mass PBHs produced from the model can constitute the whole dark matter [92, 93, 94], and is in excellent agreement with ACT observational constraints.
@@
It is also worth mentioning that, as illustrated in Fig. 4, the enhancement of the small-scale power spectrum not only affects the large-scale spectral index $n_{s}$ and its running $\alpha_{s}$, which can be probed through direct measurements of the primordial power spectrum from future CMB surveys [95, 96], but also induces significant $\mu$-distortion [97, 98]—potentially detectable by PIXIE [29, 28] [^6] See detailed analysis in [11].. Additionally, the associated induced gravitational waves fall within the sensitivity ranges of future detectors such as LISA [31], DECIGO [100], Taiji [101]/TianQin [102] and BBO [71]. Thus, upcoming CMB and gravitational wave missions may provide critical observational tests of this scenario.
Figure 5: The $n_{s}-r$ relation in the $R^{2}$ model and $\chi$-extended $R^{2}$ model shown alongside the joint constraints from BICEP/Keck observations [103] (denoted as BK18) with Planck-LB (orange), ACT-LB (blue), and P-ACT-LB (purple). The constraint contours are taken from [6]. The dark blue dashed line with hollow dots refers to the $n_{s}-r$ relation in the $R^{2}$ model, and the light cyan dashed line and markers represent the $\chi$-extended $R^{2}$ model with PBH production. For the light cyan plot, we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From lower $n_{s}$ (left) to higher $n_{s}$ (right), the coupling constants used are $\xi=0.3125$ (hollow dot), 0.07250.07250.0725, 0.07150.07150.0715, 0.07050.07050.0705, 0.06950.06950.0695, 0.06890.06890.0689 (star-shaped, Case 2), 0.06850.06850.0685.
Figure 6: The $n_{s}-\alpha_{s}$ relation in $\chi$-extended $R^{2}$ model shown alongside the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). The light cyan dots refer to the $\alpha_{s}-n_{s}$ relation in $\chi$-extended $R^{2}$ model with PBH production, for which we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), 0.07250.07250.0725, 0.07150.07150.0715, 0.07050.07050.0705, 0.06950.06950.0695, 0.06890.06890.0689 (star-shaped, Case 2), 0.06850.06850.0685.
@@ V Discussion and Conclusion
The recently released ACT data, combined with Planck observations, report a slightly larger scalar spectral index $n_{s}$ and a positive running $\alpha$ placing well-motivated $R^{2}$ inflation under tension at $2\sigma$ level. In this paper, we have shown that the tension can be resolved by introducing a non-minimally coupled scalar field $\chi$ to the $R^{2}$ model. By tuning the non-minimal coupling constant $\xi$, the model realizes a hybrid-like inflationary scenario, in which the non-minimally coupled scalar field $\chi$ contributes a blue-tilted component to the primordial power spectrum, yielding an increase in the $n_{s}$ and a small but positive $\alpha_{s}$, both consistent with current observational constraints.
At the same time, the blue-tilted power spectrum significantly enhances the fluctuation amplitude at small scales, potentially leading to the formation of primordial black holes (PBHs) that could constitute the entire dark matter abundance. In Fig. 3, we illustrate the relationship between the spectral index $n_{s}$, its running $\alpha_{s}$, and the resulting PBH mass under the assumption that PBHs account for all dark matter.
Consistency with P-ACT observations implies that a positive running ($\alpha_{s}>0$) may serve as indirect evidence for a blue-tilted spectrum and enhanced small-scale power, which can lead to the production of PBHs with masses $\lesssim 10^{20}~\mathrm{g}$ that constitute the entire dark matter content. Notably, the constraint from $\alpha_{s}-n_{s}$ allows partly the asteroid-mass window for PBH dark matter ($10^{17}\mathrm{g}\lesssim M_{\mathrm{PBH}}\lesssim 10^{23}\mathrm{g}$) and the extended mass window if the memory burden effect is considered ($10^{5}\mathrm{g}\lesssim M_{\mathrm{PBH}}\lesssim 10^{10}\mathrm{g}$ when memory-burden parameter $n_{\mathrm{MB}}=2$ [104, 105, 106]).
We emphasize that our model does not require PBHs to account for all dark matter in order to resolve the observed tension. Rather, it naturally correlates modifications to the large-scale spectrum with potentially observable small-scale features.
We also comment here on the potential observational signatures of the primordial blue-tilted power spectrum predicted by our model.
A slightly blue-tilted spectrum with a peak at small scales would be a distinctive prediction of this scenario. The blue tilt of the power spectrum can be probed by upcoming measurements of CMB spectral distortions, such as the $\mu$-distortion detectable by experiments like PIXIE. Meanwhile, the associated peak in the power spectrum can generate a second-order gravitational wave background, which may be within the reach of future detectors such as LISA [31], DECIGO [100], Taiji [101]/TianQin [102] and BBO [71].
In addition, recent studies suggest that a blue-tilted primordial power spectrum at $k\gtrsim 1~\mathrm{Mpc}^{-1}$ could lead to the early formation of high-redshift dark matter halos [107, 108, 109], and enhance the substructure abundance within low-redshift dark matter halos [110]. These features may help to explain the unexpectedly large number of massive galaxies observed at high redshift by the James Webb Space Telescope (JWST) [111, 112], the ’too many satellite problem’ [113, 114] as well as the anomalous flux ratio problem [115] detected in strong gravitational lensing systems.
Finally, we should stress the intriguing connection between the inflation scale and the small neutrino masses $m_{\nu_{i}}$. The $\chi$ boson has coupling to the right handed neutrinos as $g_{i}\chi N_{Ri}N_{Ri}$ which generates the large Majorana masses $M_{i}$ for the right-handed neutrinos after the total inflation ends as explained in the introduction. Through the seesaw mechanism [38, 39, 40, 41], the observed neutrino masses $m_{\nu_{i}}$ is given by $m_{\nu_{i}}\simeq h_{i}^{2}\langle H\rangle^{2}/M_{i}$. Here, the $H$ is the standard-model Higgs boson, $M_{i}\simeq g_{i}\langle\chi\rangle$, and the $\langle\chi\rangle=v$ is determined by the inflation scale as shown in the text. Therefore, it is very interesting that there is a strong correlation between the inflation scale and the observed small neutrino masses.
@@ Acknowledgements.
We thank Misao Sasaki and Ying-li Zhang for useful discussions. This work is supported in part by JSPS KAKENHI Grant Nos. JP20H05853, JP24K00624 [X.W.]; JP23KF0289, JP24H01825, JP24K07027 [K.K.]; No. JP24H02244 [T.T.Y.], and by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo [X.W.]. Kavli IPMU is supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. X.W. would also like to acknowledge the insightful input from the ’Random Discussion’ and the ’Workshop on the Physics and Mathematics of the Universe’ held at Kavli IPMU, which greatly contributed to the completion of this work.